Prove that for every natural number


n, we can express nk, where k≥ 2 is an integer, as a sum of n consective odd natural numbers.

Grade:10

1 Answers

Radhika Batra
14 years ago

Dear Ajinkya,

you have to prove that 

1+3+5+7+ -------- +n  = nk   ;while k>= 2.

Therefore : 1+3+5+7 ----------+ n  = [2a +(n-1)d]n/2 ; a= first term; d = common difference.

put the value in the formulae

thus   1+3+5+7+ ------------+ n = [2*1 +(n-1)*2]n/2 = n2   which is true for k>= 2 

askIITians.com provides online iit jee courses and IIT JEE Test Series with IITians. Click here to get free online test series and check your status timely or you can join us as our registered user for getting best iit jee study material or iit jee test series.

Click here to find IIT JEE Packages or course details with askIITians.

for further details visit the folowing links:

http://www.askiitians.com//iit-jee-online-coaching.aspx

http://www.askiitians.com/iit-jee-test-series.aspx

http://www.askiitians.com/iit-study-material/

Askiitians provide you all the best possible information, feel free to ask.

all the best.

thanks and regards.

Akhilesh Shukla